Please help:
Show that the equation 6 cos theta + 5 tan theta = 0
may be rewritten in the form 6 sin^2 theta - 5 sin theta - 6 = 0

Answer by Alan3354(69443)

Show that the equation 6 cos theta + 5 tan theta = 0
tan = sin/cos
Multiply thru by cos
6cos^2 + 5sin = 0
cos^2 = 1 - sin^2
6(1 - sin^2) + 5sin = 0
6(sin^2 - 1) - 5sin = 0
6sin^2 - 6 - 5sin = 0
--------------
may be rewritten in the form 6 sin^2 theta - 5 sin theta - 6 = 0